OpenStudy (wintersuntime):

Rider A: y=3/2x Rider B: y=1/2x+10 Use the equal values method to determine what time rider A passed RiderB: How far from the starting line were the two racers at that time ?

OpenStudy (wintersuntime):

@mathmale

OpenStudy (wintersuntime):

@kittiwitti1

OpenStudy (kittiwitti1):

Si?

OpenStudy (wintersuntime):

I need help with that question

OpenStudy (wintersuntime):

I need help solving it

OpenStudy (kittiwitti1):

I don't know e_e

OpenStudy (kittiwitti1):

I'm sorry ; -;

OpenStudy (wintersuntime):

2 (3/2x=1/2x+10)

OpenStudy (kittiwitti1):

! Oh wait... you set the functions equal to each other right

OpenStudy (wintersuntime):

yes

OpenStudy (kittiwitti1):

Oh, okay. Yeah you're on the right path (:

OpenStudy (wintersuntime):

Then I distribute ?

OpenStudy (kittiwitti1):

Yep ☺

OpenStudy (wintersuntime):



OpenStudy (kittiwitti1):

\[2(\frac{3}{2}x)=2(\frac{1}{2})+2(10)\rightarrow 3x=1+20\]So 3x=21

OpenStudy (kittiwitti1):

Yep ♪

OpenStudy (wintersuntime):

Then add 3x on both sides

OpenStudy (kittiwitti1):

Whoa, going a little overboard there lol Nope, did you see what I did ♦

OpenStudy (wintersuntime):

subtract 3x on both sides lol

OpenStudy (kittiwitti1):

Um.. No o_o?

OpenStudy (wintersuntime):

divide

OpenStudy (kittiwitti1):

Yeah ☺

OpenStudy (wintersuntime):

x=7

OpenStudy (kittiwitti1):

Yes!! ☺☺

OpenStudy (wintersuntime):

I don't know on what equation I should plug it in

OpenStudy (kittiwitti1):

Either one works I think

OpenStudy (wintersuntime):

So I could plug it in with 1/2x+10 right ?

OpenStudy (wintersuntime):

Is that wrong or right ??

OpenStudy (kittiwitti1):

Sorry! Yeah, it works ♦

OpenStudy (wintersuntime):

its okay

OpenStudy (wintersuntime):

I got 13.5

OpenStudy (kittiwitti1):

??? That's right but it's weird o_e

OpenStudy (wintersuntime):

or as a fraction is 27/2

OpenStudy (kittiwitti1):

OH WE MESSED UP

OpenStudy (kittiwitti1):

It's x+20 not 1+20

OpenStudy (wintersuntime):

Wait for the equation ??

OpenStudy (kittiwitti1):

Yeah lol I'm so sorry > <

OpenStudy (wintersuntime):

its okay

OpenStudy (wintersuntime):

So all of its wrong?

OpenStudy (kittiwitti1):

Okay so it's \[\frac{1}{2}x=\frac{3}{2}x+10\]\[2\left(\frac{1}{2}x\right)=2\left(\frac{3}{2}x\right)+2(10)\]

OpenStudy (kittiwitti1):

It's not wrong, we just left out an x. I'm redoing it so we don't mess up again is all ♦

OpenStudy (kittiwitti1):

\[\rightarrow x=3x+20\]\[-4x=20...\]

OpenStudy (wintersuntime):

x=-5

OpenStudy (wintersuntime):

Is it right ?

OpenStudy (kittiwitti1):

Yes ☺

OpenStudy (wintersuntime):

plug it in and y=2.5 or as a fraction 5/2

OpenStudy (kittiwitti1):

Yes ♪

OpenStudy (kittiwitti1):

Except it's negative, don't forget that

OpenStudy (wintersuntime):

oh okay

OpenStudy (wintersuntime):

thanks

OpenStudy (kittiwitti1):

np ☺